# Exploring Self-Gravitating Cylindrical Structures in Modified Gravity: Insights from Scalar-Vector-Tensor Theory

### v2

@@ Abstract
We investigate static cylindrical solutions within an extended theory of modified gravity. By incorporating various coupling functions through a straightforward boost symmetry approach, we establish the equations of motion in a self-consistent manner and subsequently determine the linear scalar field profile. Utilizing analytical methods, we solve the system of equations for the metric functions and the $U(1)$ gauge field, revealing their dependence on Bessel’s functions. To comprehend gravito-objects exhibiting cylindrical symmetry, we develop a perturbative framework aimed at identifying all nontrivial solutions for the scalar profiles. Introducing first-order truncated perturbation equations for the gauge field, synchronized with metric gauges and electromagnetic field considerations, we demonstrate their integrability and obtain solutions through quadrature. Our findings suggest the feasibility of obtaining self-gravitating cylindrical structures within the scalar-vector-tensor theory.
@@ I Introduction
While Einstein’s theory of gravity remains highly successful in explaining gravitational phenomena, the need to extend general relativity arises to address various challenges, including the cosmological constant problem, dark energy, dark matter, and the quest for a quantization scheme. One straightforward extension involves introducing a scalar field, known as the scaleron degree of freedom. However, such modifications may introduce higher-order time derivatives in the equations of motion, potentially leading to Ostrogradski instabilities [1].
@@
In certain cases, these higher-order time derivatives can be formally transformed into standard second-order forms, as demonstrated by Horndeski [2],[3]. He showed that the most general second-order, covariant scalar-tensor theories lead to stable Horndeski interactions with specific coupling functions [3]. Additionally, the inclusion of an auxiliary vector field in modified theories, meaningfully coupled to other aspects of the gravitational sector, leads to the most general vector-tensor theories with second-order equations of motion. These extensions, following Horndeski’s terminology, result in specific viable couplings between the vector field and the curvature tensor, reminiscent of Maxwell’s equations under carefully chosen gauge fixing schemes. Relaxing these gauge fixings leads to a broader class of theories known as generalized Proca theories.
@@
Further developments have identified novel purely intrinsic vector interactions without scalar counterparts, such as non-minimally coupled theories to the double dual Riemann tensor, with significant cosmological and astrophysical implications. The scalar-vector-tensor (SVT) theory emerges as a unification of Horndeski scalar-tensor theories and generalized Proca theories, motivated by recent works [4]. This class of modified gravity theories presents new solutions for hairy black holes and offers alternative scenarios for dark energy [5]-[36].
@@
In this paper, we focus on investigating static, non-rotating cylindrical solutions within the SVT framework. The paper is organized as follows: Section 2 provides a brief overview of the modified scalar-vector-tensor theories of gravity. In Section 3, we propose a static Weyl’s gauge suitable for describing cylindrical objects. Section 4 is dedicated to deriving equations of motion for all fields, while Section 5 presents exact solutions for linear and quadratic scalaeron profiles. Section 6 discusses the integrability of perturbative equations of motion in detail, and the final section offers concluding remarks.
@@ II Heisenberg’s Gauge-invariant scalar-vector-tensor theories
The objective of this section is to provide a foundational framework for understanding the Lagrangian of the innovative modified gravity theories recently proposed in Ref. [1]. Referred to as scalar-vector-tensor gravity theories, this class of models is widely regarded as a fusion of Horndeski theories and generalized Proca theories. The SVT framework exhibits gauge dependency and necessitates meticulous gauge fixing. By adopting a suitable gauge-invariant platform, one arrives at the model introduced in Ref.[1]. The Lagrangian densities of typical scalar-vector-tensor theories, incorporating arbitrary interaction terms (coupling functions), can be represented as follows:
@@
The renowned SVT framework offers a diverse landscape for formulating and exploring theories that extend beyond scalar-vector-tensor formulations. This can be achieved by implementing disformal transformations [36]. A plausible and physically meaningful interpretation of the interactions in ${\cal L}^{2,3,4}_{\rm SVT}$ is to view them as generators of disformal transformations or coupling functions [37].
@@ III Cylindrical geometry
Adhering to this symmetry class, we assume all metric functions to be solely functions of $r$, denoted as $U=U(r)$ and $K=K(r)$. Exact vacuum solutions to the Einstein field equations with this symmetry can be attained, leading to the discovery of the Levi-Civita solution. Extensive investigations into cylindrical solutions across different gravitational theories have been conducted in the past, primarily aimed at understanding topological defects through Riemannian geometry [12]. Inspired by string theory, this approach elucidates how topological defects mimic cylindrical solutions.
In general relativity (GR), the simplest cylindrical model, presented as an exact class of metrics, was discovered by Kasner in Refs. [13]. Linet subsequently proposed the first extension of the Kasner solution in an exact form [14]. Tian later generalized the Linet solution by incorporating a cosmological constant [15], offering the capability to study both cosmological and AdS limits in cylindrical systems. Within the framework of modified gravity theories, the first extension of the GR solution with cylindrical symmetry in $f(R)$ gravity was presented in [16]. Matter contents can be weakly coupled to the gravitational sector without violating the equivalence principle, as studied as a source for cosmic strings by Harko in Ref. [17].
Given the origin of these solutions in string theory, higher-order corrections to the GR solutions have garnered importance, such as the Gauss-Bonnet corrections to cosmic strings investigated in [18]. When Lorentz symmetry is broken in high-energy regimes, non-relativistic theories emerge as alternatives to GR, modifying cylindrical geometries [19]. Harko and Lake made a remarkable observation, discovering a naive relation between the geometries of cosmic strings and the Bose-Einstein condensed phase of matter [20]. Further corrections to GR, such as mimetic complements to GR, also support the existence of cylindrical gravitational objects [21]. In teleparallel theories, where gravity is viewed as the effect of torsion rather than curvature, it is also feasible to find metrics for cosmic strings [22]. Moreover, cylindrical solutions have been studied in various contexts, including brane-world scenarios [23], Kaluza-Klein models [24], Lovelock Lagrangians [25], Born-Infeld [26], bimetric theories [27], scalar-tensor theories [28], Brans-Dicke theory [29], and dilation gravity [30].
As indicated in Ref. [1], the SVT theories offer the possibility of accommodating new black object solutions, marking one of the pivotal applications of these theories. To commence our inquiry, we will first delve into the cylindrical solutions within the framework of these theories. To explore these potential solutions, let us examine the following Lagrangian, with a metric signature of (+,,,)(+,-,-,-)( + , - , - , - ):
@@ IV Equations of motion
In the preceding section, we derived the reduced point-like Lagrangian for the cylindrical gravitational objects, as presented in Eq. (17). This reduced Lagrangian comprises four unknown functions denoted as $q^{a}={U,K,\pi,A}$. In the language of dynamics, obtaining the equations of motion (EoMs) is straightforward using the standard Euler-Lagrange equations, which are expressed as follows:
@@
$\boxed{U+2-4(U^{\prime}+rU^{\prime\prime})=\sum_{n=1}^{5}h_{n}(U,K,A^{\prime},A^{\prime\prime})(\pi^{\prime\prime})^{n}},$ (22)
@@
$\boxed{K-4\beta\,\left(A^{\prime}\right)^{2}{{\rm e}^{-2\,K}}-2=\sum_{n=1,3,5}p_{n}(U,K,U^{\prime},A^{\prime})(\pi^{\prime\prime})^{n}}\,,$ (28)
@@
$\boxed{A-4\,\beta\,{{\rm e}^{-2\,K}}{\it A^{\prime\prime}}+8\,{{\rm e}^{-2\,K}}\beta\,{\it A^{\prime}}{\it K^{\prime}}=\sum_{n=1}^{5}m_{n}(U,K,U^{\prime},K^{\prime},U^{\prime\prime})(\pi^{\prime\prime})^{n}}\,,$ (37)
@@
$m_{4}(U,K,U^{\prime},K^{\prime},U^{\prime\prime})={{\rm e}^{4\,U-6\,K}}V{{}_{3}}\left(\pi\right){\it U^{\prime}}\,,$ (41)
@@ V.1 Linear $\pi$ profile
Taking the derivative of Eq. (48) with respect to $r$, we obtain:
@@ V.2 Solutions with dilaton quadratic profile
$\boxed{U+2-4(U^{\prime}+rU^{\prime\prime})=\sum_{n=1}^{5}h_{n}(U,K,A^{\prime},A^{\prime\prime})},$ (76)
@@
$\boxed{K-4\beta\,\left(A^{\prime}\right)^{2}{{\rm e}^{-2\,K}}-2=\sum_{n=1,3,5}p_{n}(U,K,U^{\prime},A^{\prime})}\,,$ (82)
@@
$\boxed{A-4\,\beta\,{{\rm e}^{-2\,K}}{\it A^{\prime\prime}}+8\,{{\rm e}^{-2\,K}}\beta\,{\it A^{\prime}}{\it K^{\prime}}=\sum_{n=1}^{5}m_{n}(U,K,U^{\prime},K^{\prime},U^{\prime\prime})}$ (90)
@@
$\boxed{K_{1}-4\beta e^{-2K_{0}}(-2A_{0}^{\prime}K_{1}^{\prime}+A_{1}^{\prime})=-4e^{2U_{0}-4K_{0}}U_{0}^{\prime}A^{\prime}_{0}V_{2}(\pi^{0})}.$ (107)
@@ VII Conclusions
Although our inquiry refrains from delving into the intricacies of singularity theorems, our meticulously crafted metrics in both the linear and perturbative domains unfurl conical singularities, wherein the conical factor bears a poignant relationship with the mass of the cosmic string. This tantalizing revelation hints at the prospect of delineating a form of no-go theorem within this nuanced framework. In summation, our expedition engenders a panoramic vista of the SVT theory as a compelling alternative to GR, showcasing its remarkable versatility in grappling with the manifold gravitational phenomena that adorn the cosmic tapestry.

### v3

@@ Abstract
We investigate static cylindrical solutions within an extended theory of modified gravity. By incorporating various coupling functions through a straightforward boost symmetry approach, we establish the equations of motion in a self-consistent manner and subsequently determine the linear scalar field profile. Utilizing analytical methods, we solve the system of equations for the metric functions and the $U(1)$ gauge field, revealing their dependence on Bessel’s functions. To comprehend gravito-objects exhibiting cylindrical symmetry, we develop a perturbative framework aimed at identifying all nontrivial solutions for the scalar profiles. Introducing first-order truncated perturbation equations for the gauge field, synchronized with metric gauges and electromagnetic field considerations, we demonstrate their integrability and obtain solutions through quadrature. Our findings suggest the feasibility of obtaining self-gravitating cylindrical structures within the scalar-vector-tensor theory. These cylindrical structures could provide insights into the behavior of gravitational and gauge fields in modified gravity, potentially offering new perspectives on astrophysical phenomena such as cosmic strings and cylindrical gravitational waves.
@@ I Introduction
While Einstein’s theory of gravity remains highly successful in explaining gravitational phenomena, the need to extend general relativity arises to address various challenges, including the cosmological constant problem, dark energy, dark matter, and the quest for a quantization scheme. One straightforward extension involves introducing a scalar field, known as the scaleron degree of freedom. However, such modifications may introduce higher-order time derivatives in the equations of motion, potentially leading to Ostrogradski instabilities Heisenberg:2018acv .
@@
In certain cases, these higher-order time derivatives can be formally transformed into standard second-order forms, as demonstrated by Horndeski Horndeski:1974wa , Horndeski:1976gi . He showed that the most general second-order, covariant scalar-tensor theories lead to stable Horndeski interactions with specific coupling functions Horndeski:1976gi . Additionally, the inclusion of an auxiliary vector field in modified theories, meaningfully coupled to other aspects of the gravitational sector, leads to the most general vector-tensor theories with second-order equations of motion. These extensions, following Horndeski’s terminology, result in specific viable couplings between the vector field and the curvature tensor, reminiscent of Maxwell’s equations under carefully chosen gauge fixing schemes. Relaxing these gauge fixings leads to a broader class of theories known as generalized Proca theories.
@@
Further developments have identified novel purely intrinsic vector interactions without scalar counterparts, such as non-minimally coupled theories to the double dual Riemann tensor, with significant cosmological and astrophysical implications. The scalar-vector-tensor (SVT) theory emerges as a unification of Horndeski scalar-tensor theories and generalized Proca theories, motivated by recent works Heisenberg:2014rta . This class of modified gravity theories presents new solutions for hairy black holes and offers alternative scenarios for dark energy Jimenez:2016isa -Tsujikawa:2015upa .
Previous studies have primarily focused on spherical solutions in modified gravity theories, which have provided significant insights but also revealed certain limitations. Spherical solutions are often idealized and may not capture the full complexity of various astrophysical and cosmological structures. For instance, while spherical symmetry simplifies the mathematical treatment, it may not adequately describe elongated or cylindrical objects observed in the universe, such as cosmic strings or certain types of gravitational waves. Addressing these limitations necessitates exploring solutions with different symmetries.
In this context, cylindrical solutions offer a compelling avenue for further research. Cylindrical structures can model various astrophysical phenomena and provide a richer framework for understanding gravitational interactions in modified gravity theories. Additionally, previous studies on cylindrical solutions in modified gravity have been undertaken. For a more thorough background, we direct readers to Refs. Houndjo:2012sz ; Houndjo:2013us ; Momeni:2015aea ; Azadi:2008qu . These studies underscore the importance of extending our investigations to cylindrical symmetry to fully grasp the potential implications of modified gravity theories.
@@
In this paper, we focus on investigating static, non-rotating cylindrical solutions within the SVT framework. The paper is organized as follows: Section 2 provides a brief overview of the modified scalar-vector-tensor theories of gravity. In Section 3, we propose a static Weyl’s gauge suitable for describing cylindrical objects. Section 4 is dedicated to deriving equations of motion for all fields, while Section 5 presents exact solutions for linear and quadratic scalaron profiles. Section 6 discusses the integrability of perturbative equations of motion in detail, and the final section offers concluding remarks.
@@ II Heisenberg’s Gauge-invariant scalar-vector-tensor theories
SVT theory offers a comprehensive framework to address various cosmological enigmas such as dark energy, dark matter, and the accelerated expansion of the universe. One of the primary advantages of SVT theory over other modified gravity theories is its ability to naturally incorporate both scalar and vector degrees of freedom, leading to richer and more versatile interactions. This versatility allows the SVT theory to simultaneously account for the large-scale structure of the universe and local gravitational phenomena without introducing instabilities. Additionally, SVT theory can produce viable cosmological models that align with observational data, offering alternative explanations for the phenomena typically attributed to dark energy and dark matter. By unifying aspects of Horndeski scalar-tensor theories and generalized Proca theories, SVT theory provides a robust platform for exploring new gravitational dynamics and potentially resolving long-standing cosmological puzzles.
The objective of this section is to provide a foundational framework for understanding the Lagrangian of the innovative modified gravity theories recently proposed in Ref. [1]. Referred to as scalar-vector-tensor gravity theories, this class of models is widely regarded as a fusion of Horndeski theories and generalized Proca theories. The SVT framework exhibits gauge dependency and necessitates meticulous gauge fixing. By adopting a suitable gauge-invariant platform, one arrives at the model introduced in Ref.Heisenberg:2018acv . The Lagrangian densities of typical scalar-vector-tensor theories, incorporating arbitrary interaction terms (coupling functions), can be represented as follows:
@@
The renowned SVT framework offers a diverse landscape for formulating and exploring theories that extend beyond scalar-vector-tensor formulations. This can be achieved by implementing disformal transformations Tsujikawa:2015upa . A plausible and physically meaningful interpretation of the interactions in ${\cal L}^{2,3,4}_{\rm SVT}$ is to view them as generators of disformal transformations or coupling functions disformal .
@@ III Cylindrical geometry
Adhering to this symmetry class, we assume all metric functions to be solely functions of $r$, denoted as $U=U(r)$ and $K=K(r)$. Exact vacuum solutions to the Einstein field equations with this symmetry can be attained, leading to the discovery of the Levi-Civita solution. Extensive investigations into cylindrical solutions across different gravitational theories have been conducted in the past, primarily aimed at understanding topological defects through Riemannian geometry vilenkin . Inspired by string theory, this approach elucidates how topological defects mimic cylindrical solutions.
In general relativity (GR), the simplest cylindrical model, presented as an exact class of metrics, was discovered by Kasner in Refs. Kasner . Linet subsequently proposed the first extension of the Kasner solution in an exact form Linet . Tian later generalized the Linet solution by incorporating a cosmological constant tian , offering the capability to study both cosmological and AdS limits in cylindrical systems. Within the framework of modified gravity theories, the first extension of the GR solution with cylindrical symmetry in $f(R)$ gravity was presented in fr . Matter contents can be weakly coupled to the gravitational sector without violating the equivalence principle, as studied as a source for cosmic strings by Harko in Ref. frlm .
Given the origin of these solutions in string theory, higher-order corrections to the GR solutions have garnered importance, such as the Gauss-Bonnet corrections to cosmic strings investigated in gb . When Lorentz symmetry is broken in high-energy regimes, non-relativistic theories emerge as alternatives to GR, modifying cylindrical geometries nonrelativistic . Harko and Lake made a remarkable observation, discovering a naive relation between the geometries of cosmic strings and the Bose-Einstein condensed phase of matter be . Further corrections to GR, such as mimetic complements to GR, also support the existence of cylindrical gravitational objects mimetic . In teleparallel theories, where gravity is viewed as the effect of torsion rather than curvature, it is also feasible to find metrics for cosmic strings teleparallel . Moreover, cylindrical solutions have been studied in various contexts, including brane-world scenarios brane , Kaluza-Klein models kk , Lovelock Lagrangians love , Born-Infeld bi , bimetric theories bit , scalar-tensor theories st , Brans-Dicke theory bd , and dilation gravity dilaton .
As indicated in Ref. Heisenberg:2018acv , the SVT theories offer the possibility of accommodating new black object solutions, marking one of the pivotal applications of these theories. To commence our inquiry, we will first delve into the cylindrical solutions within the framework of these theories. To explore these potential solutions, let us examine the following Lagrangian, with a metric signature of (+,,,)(+,-,-,-)( + , - , - , - ):
@@ IV Equations of motion
In the preceding section, we derived the reduced point-like Lagrangian for the cylindrical gravitational objects, as presented in Eq. (17). To ensure clarity and rigor in deriving the equations of motion, we follow a systematic approach that integrates the boost symmetry principles. This approach not only simplifies the form of our metric and gauge field components but also underscores their physical relevance in the context of cylindrical gravitational structures. By enforcing boost symmetry, we constrain the metric functions $U$, $K$, and $\pi$ to be functions solely of the radial coordinate $r$, while allowing the gauge field $A$ to retain dependence on both $r$ and the angular coordinate $\varphi$. This reduction in degrees of freedom enhances the tractability of our equations, facilitating a more insightful exploration of their physical implications within the framework of SVT theory. This reduced Lagrangian comprises four unknown functions denoted as $q^{a}={U,K,\pi,A}$. In the language of dynamics, obtaining the equations of motion (EoMs) is straightforward using the standard Euler-Lagrange equations, which are expressed as follows:
@@
$U+2-4(U^{\prime}+rU^{\prime\prime})=\sum_{n=1}^{5}h_{n}(U,K,A^{\prime},A^{\prime\prime})(\pi^{\prime\prime})^{n},$ (22)
@@
$K-4\beta\,\left(A^{\prime}\right)^{2}{{\rm e}^{-2\,K}}-2=\sum_{n=1,3,5}p_{n}(U,K,U^{\prime},A^{\prime})(\pi^{\prime\prime})^{n}\,,$ (28)
@@
$A-4\,\beta\,{{\rm e}^{-2\,K}}{\it A^{\prime\prime}}+8\,{{\rm e}^{-2\,K}}\beta\,{\it A^{\prime}}{\it K^{\prime}}=\sum_{n=1}^{5}m_{n}(U,K,U^{\prime},K^{\prime},U^{\prime\prime})(\pi^{\prime\prime})^{n}\,,$ (37)
@@
$m_{4}(U,K,U^{\prime},K^{\prime},U^{\prime\prime})={{\rm e}^{4\,U-6\,K}}V{\it{}_{3}}\left(\pi\right){\it U^{\prime}}\,,$ (41)
@@ V.1 Linear $\pi$ profile
The utilization of Bessel’s functions in our analysis stems from their intrinsic properties that align with cylindrical symmetry and radial dependence in the gravitational field equations. Specifically, Bessel functions naturally arise in scenarios where the underlying geometry exhibits cylindrical symmetry, such as in our study of self-gravitating cylindrical structures within modified gravity theories. These functions satisfy differential equations that emerge from the cylindrical coordinates and boundary conditions typical in such gravitational configurations. Moreover, their orthogonality properties and well-defined behavior under varying boundary conditions make them indispensable tools for solving partial differential equations in cylindrical geometries, thereby providing a robust mathematical framework to explore the structural and dynamical aspects of our scalar-vector-tensor theory. Taking the derivative of Eq. (48) with respect to $r$, we obtain:
@@ V.2 Solutions with dilaton quadratic profile
$U+2-4(U^{\prime}+rU^{\prime\prime})=\sum_{n=1}^{5}h_{n}(U,K,A^{\prime},A^{\prime\prime}),$ (76)
@@
$K-4\beta\,\left(A^{\prime}\right)^{2}{{\rm e}^{-2\,K}}-2=\sum_{n=1,3,5}p_{n}(U,K,U^{\prime},A^{\prime})\,,$ (82)
@@
$A-4\,\beta\,{{\rm e}^{-2\,K}}{\it A^{\prime\prime}}+8\,{{\rm e}^{-2\,K}}\beta\,{\it A^{\prime}}{\it K^{\prime}}=\sum_{n=1}^{5}m_{n}(U,K,U^{\prime},K^{\prime},U^{\prime\prime})$ (90)
@@
The perturbative approach employed in our study plays a pivotal role in analyzing the stability and behavior of cylindrical solutions within the scalar-vector-tensor (SVT) theory. We adopt standard perturbation techniques, leveraging Maple software for computational assistance. The method involves expanding the metric and field variables around known solutions in a series of small parameter perturbations. This allows us to systematically explore deviations from exact solutions, providing insights into the stability and physical implications of our cylindrical configurations. By breaking down the perturbation process into manageable steps and utilizing computational tools, we aim to elucidate how slight modifications in the scalar and vector fields affect the overall gravitational and electromagnetic profiles of these cylindrical structures.
@@
$K_{1}-4\beta e^{-2K_{0}}(-2A_{0}^{\prime}K_{1}^{\prime}+A_{1}^{\prime})=-4e^{2U_{0}-4K_{0}}U_{0}^{\prime}A^{\prime}_{0}V_{2}(\pi^{0}).$ (107)
@@
The first-order truncated perturbation equations derived in our study offer valuable insights into the physical implications of cylindrical solutions within the scalar-vector-tensor (SVT) theory. These equations provide a refined understanding of how small deviations from exact solutions manifest in the gravitational and electromagnetic fields surrounding these cylindrical structures. By solving these equations, we uncover the intricate interplay between the scalar and vector fields, shedding light on the stability and observable characteristics of these gravito-objects. Furthermore, these solutions contribute to theoretical predictions, offering potential explanations for observational phenomena such as gravitational lensing effects or deviations from standard gravitational wave signatures. This analytical framework not only advances our theoretical understanding but also provides a basis for future empirical tests and observations in astrophysical contexts.
@@ VII Conclusions
In addition to establishing the stability and perturbative behavior of cylindrical solutions within SVT theory, it is pertinent to note the implications of conical singularities and the theoretical considerations regarding the mass of cosmic strings. These aspects, while touched upon in the broader context of modified gravity theories, represent avenues for further investigation beyond the scope of this study. Understanding the implications of conical singularities and the dynamics of cosmic strings within SVT theory not only enriches our theoretical understanding but also offers potential insights into their observable effects in astrophysical and cosmological contexts.
Although our inquiry refrains from delving into the intricacies of singularity theorems, our meticulously crafted metrics in both the linear and perturbative domains unfurl conical singularities, wherein the conical factor bears a poignant relationship with the mass of the cosmic string. This tantalizing revelation hints at the prospect of delineating a form of no-go theorem within this nuanced framework. In summation, our expedition engenders a panoramic vista of the SVT theory as a compelling alternative to GR, showcasing its remarkable versatility in grappling with the manifold gravitational phenomena that adorn the cosmic tapestry. Further phenomenological studies represent an intriguing future direction for this research. For instance, investigating the impact of cylindrical structures on the large-scale universe structure holds significant potential. Moreover, exploring the formation mechanisms and stability criteria of cosmic strings within the SVT theory will shed light on the conditions under which these structures can form and endure.